# Taylor’s hypothesis and its impact on flux measurements in a forest clearcut

### v1

@@ 2.3 DTS data processing
DTS data were post-field calibrated with Python (3.11.9) scripts and the processing followed des Tombe et al. (2020), Van De Giesen et al. (2012) and Peltola et al. (2022) with slight differences. Temperature can be estimated from the observed Stokes ($P_{s}$) and anti-Stokes ($P_{as}$) signals with the following equation (Van De Giesen et al., 2012):
[…]
@@
$\sigma_{T}^{2}(s)=\frac{\overline{\left(T(s,t)-T(s+\Delta s,t)\right)^{2}}}{2},$ (2)
where $\overline{()}$ denotes temporal averaging and $\Delta s$ is the spatial step along the cable (0.127 cm). In this equation, we assume that the variability between consecutive data points in space is dominated by noise. The calibrated temperature $T$ at each $t$ and $s$ was calculated as a weighted average of $T_{f}$ and $T_{r}$ where inverse of noise estimates were used as weights (des Tombe et al., 2020). More importantly, the noise estimates for $T$ were calculated following noise propagation of weighted averaging. The spatial resolution of the data were reduced to 30 cm with spatial aggregation using weighted averaging as above.
After this calibration procedure, we obtained calibrated $T$ at each time step $t$ and location $s$ along the cable. The $T$ data measured along the cable were converted to physical coordinates (ETRS89/TM35FIN coordinate system) using the measured cable locations (Sect. 2.2). Since there were double cabling at each physical location, the double measurements made at each location were further averaged using weighted averaging as above. By utilizing this procedure, we obtained calibrated $T$ data with 2.3 s temporal and 30 cm spatial resolution throughout the measurement campaign at well-defined physical ($x$,$y$,$z$) locations (red lines in Fig. 1) over the study area. These calibrated temperature values were compared against the sonic anemometer measurements at the clearcut showing satisfactory agreement between the two (not shown).
@@ 2.4 Convection speed estimation
Figure 2: (a) The space-time contour plot of temperature fluctuations ($T^{\prime}$) is shown from the DTS measurements of a specific section (Section-1) for a particular 30-min period between 10:30-11:00 UTC on 18 May 2024. The spatial locations along the cable section are denoted by $X$ and the temperature fluctuations are obtained by linearly detrending the temperature signals measured along time. The white horizontal strips indicate the regions where the DTS cable was fastened to masts and the temperature values were omitted from those locations. (b) The space-time correlation curves of temperature fluctuations ($R_{TT}(r,\tau)$) are shown for the same 30-min period obtained from Section-1. The temporal lags or leads are denoted as $\tau$ while the colors indicate different values of spatial separations along $X$, $r$. The blue squares denote the peaks of the correlation curves ($\tau_{p}$) for each value of $r$. (c) A scatter plot between $r$ and $\tau_{p}$ is shown. The thick black color denotes the best fit line to compute the convective speed of the temperature structures.
@@
To use the DTS and EC datasets, we focus on buoyancy-dominated conditions when the temperature signals were sufficiently strong. In particular, we selected those 30-min periods when there was no rain, incoming shortwave radiation was larger than 300 $\rm W\ m^{-2}$ and it was not strongly varying within the 30-min periods, the sensible heat flux ($>50\rm\ Wm^{-2}$) and variance of temperature ($>0.1\rm\ K^{2}$) were sufficiently large, air temperature did not exhibit strong trends, and the direction of the mean wind was within ±$\pm 20^{\circ}$ of the DTS cable positions. This wind direction criterion resulted in two different scenarios. For Sections 1 to 5, the mean wind was parallel to the forest edge while for Section 6 it was perpendicular to the edge (see Fig. LABEL:fig:1b). After applying all these conditions, we ended up with a total of 35 hours of DTS and EC data, or in other words, 70 half-hour periods. The DTS dataset was used to estimate the convective speeds of the turbulent structures by using the temperature signal as a tracer. The EC dataset served as a testbed to evaluate the impact of these convective speeds on the routine flux measurements.
@@
Out of all our selected runs, for a specific representative 30-min period (10:30-11:00 UTC, 18 May 2024), we show the space-time contour plot of temperature fluctuations from Section 1 of the DTS cable (Fig. 2a). The temperature values were omitted from those locations where the DTS cable was fastened to wooden structures. Nevertheless, to ensure continuity in the data, those missing values were replaced with linearly-interpolated ones. Thereafter, the temperature fluctuations ($T^{\prime}$) were computed by linearly detrending the data over time separately at each location along the cable sections. The distance along Section 1 is denoted as $X$. To compute the convective speed of the temperature structures, we construct the space-time correlation curves from the DTS dataset. Mathematically, the two-point correlation functions of temperature fluctuations can be defined as,
@@
where $\tau$ are the temporal lags or leads, $r$ denotes the spatial separations along $X$, and $\sigma_{T(X)}$, $\sigma_{T(X+r)}$ are the standard deviations of temporal temperature fluctuations at locations $X$ and $X+r$. For the same data as shown in Fig. 2a, the quantity $R_{TT}(\tau,r)$ are plotted in Fig. 2b. Here, different colors represent the increasing values of $r$ while the blue squares indicate the peaks ($\tau_{p}$) attained by $R_{TT}(\tau,r)$ for a specific $r$. By fitting a straight line to a scatter plot between $r$ and $\tau_{p}$, the bulk convective speed of the temperature structures can be estimated (see the solid black line in Fig. 2c). The goodness of the fit can be adjudged by its $R^{2}$ value, which was well-beyond 0.9 in this example case (Fig. 2c). This linear relationship between $\tau_{p}$ and $r$ can be described as, $r=U_{c}\tau_{p}$, where $U_{C}$ is the bulk convective speed ($\rm m\ s^{-1}$). We restrict the $r$ values within 21 m while making the scatter plot, since for $r>21\rm\ m$, the $R_{TT}(\tau,r)$ values are small and therefore the estimations of $\tau_{p}$ are not statistically robust.
To test the robustness of this method for atmospheric surface layer flows, we repeated this analysis on the Cheng et al. (2017) dataset and the results yielded similar features (see Fig. S1 of the Supplementary Material). However, this particular method does not distinguish between positive and negative structures, or the possible scale-dependence of the convection velocities. In convective flows, a significant asymmetry is observed between the positive and negative temperature structures, which carry the signatures of the buoyant updrafts and downdrafts (Chowdhuri et al., 2020b). This asymmetry is related to how the flow is organized under such conditions. We illustrate this aspect in Fig. 9 of Appendix A. Therefore, it is prudent to investigate whether these two structures move at similar speeds. In the absence of spatial vertical velocity measurements, we associate these updrafts and downdrafts with positive and negative temperature fluctuations, and refer to them as up- and down-plumes, respectively.
Figure 3: The contours of joint probability density function (JPDF) of temperature fluctuations separated by a spatial distance $r$ are shown in subplot (a) for the same period as used in Fig. 2. The contour values are represented by a colorbar while the different quadrants signify how the signs of the temperature fluctuations co-vary. The thick dash-dotted black line represents a straight line with a slope equal to the Pearson correlation coefficient between the two temperature signals. The space-time correlation curves are shown for two different scenarios, by conditioning the temperature fluctuations based on (b) when they are only positive, i.e. relate to the quadrant I in subplot (a) ($R^{+}_{TT}(r,\tau)$) or (c) when they are only negative, i.e. relate to the quadrant III ($R^{-}_{TT}(r,\tau)$). The colors indicate different values of $r$ and the pink and blue squares denote the peaks of the conditional correlation curves ($\tau^{+}_{p}$ and $\tau^{-}_{p}$) for each value of $r$. (d) A scatter plot between $r$ and the peak positions of the conditional correlation curves is shown. The solid pink and blue colors denote the best fit lines to compute the convection speeds of the positive and negative temperature structures.
To carry out that exercise, we first look at the joint probability density functions (JPDFs) of temperature fluctuations at two distinct spatial locations along Section 1. These are denoted as, $T^{\prime}(X+r)$ and $T^{\prime}(X)$ and subsequently normalized by their respective standard deviations. From Fig. 3a it is clear that the JPDFs between $T^{\prime}(X+r)/\sigma_{T_{X+r}}$ and $T^{\prime}(X)/\sigma_{T_{X}}$ is skewed towards the third quadrant (III), where $T^{\prime}(X+r)<0,T^{\prime}(X)<0$. In other words, there exists a bias between the quadrants I ($T^{\prime}(X+r)>0,T^{\prime}(X)>0$) and III, which potentially could cause differences between the convective speeds of positive and negative temperature values. To lend credence to this hypothesis, we construct the conditional space-time correlation curves where $R_{TT}(r,\tau)$ values are computed based on whether $T^{\prime}(X+r)$ and $T^{\prime}(X)$ lie in quadrant I or III of Fig. 3a. This can be expressed as,
$R^{+}_{TT}(\tau,r)=\frac{\overline{T_{\rm I}^{\prime}(X,t)T_{\rm I}^{\prime}(X+r,t+\tau)}}{\sigma_{T_{\rm I}(X)}\sigma_{T_{\rm I}(X+r)}},R^{-}_{TT}(\tau,r)=\frac{\overline{T_{\rm III}^{\prime}(X,t)T_{\rm III}^{\prime}(X+r,t+\tau)}}{\sigma_{T_{\rm III}(X)}\sigma_{T_{\rm III}(X+r)}}.$ (4)
The subscripts I and III in Eq. 4 indicate that the temperature fluctuations are either sampled from the positive or negative quadrants. In Figs. 3b–c, $R^{+}_{TT}(\tau,r)$ and $R^{-}_{TT}(\tau,r)$ curves are shown for different values of $r$ and $\tau$. The pink and blue squares denote the peaks of $R^{+}_{TT}(\tau,r)$ and $R^{-}_{TT}(\tau,r)$ curves at a specific $r$ value. These peak positions are represented as $\tau^{+}_{p}$ and $\tau^{-}_{p}$, respectively. Similar to Fig. 2c, the straight line fits to the scatter plots of $r$ and $\tau^{\pm}_{p}$ (with $R^{2}>0.9$) are used to compute the sign-dependent convection speeds, expressed as, $U^{+}_{c}$ and $U^{-}_{c}$ respectively (Fig. 3d).
The same procedure was repeated for the other sections of the DTS cable to obtain $U_{c}$, $U^{+}_{c}$, and $U^{-}_{c}$. Specific to cable Section 6, we ignored the portion of the cable that resided within the forest. This was done to alleviate any additional complexities associated with sub-canopy scale eddy motions.
@@ 2.5 Scale-dependent convection speed
To compute the scale-dependent convection speeds of temperature structures, we follow the same procedure as outlined in Cheng et al. (2017). By borrowing the concept of phase-dependent convection speeds from Del Alamo and Jiménez (2009), Cheng et al. (2017) defined the convection speed at a wavenumber $\kappa$ ($U_{c}(\kappa)$) as,
$U_{c}(\kappa)=\frac{\Delta\varphi}{\kappa\times\Delta t}$ (5)
Here, $\Delta\varphi$ indicates the Fourier phase differences between the spatial temperature signals at time instants $t$ and $t+\Delta t$, $\kappa$ is the spatial wavenumber directly evaluated from the DTS measurements ($\kappa=2\pi/\lambda$, where $\lambda$ is the wavelength), and $\Delta t$ is the temporal lag (2.3 s for our DTS data). Unlike Cheng et al. (2017), we used a wavelet based approach (Torrence and Compo, 1998) to compute the phase differences between the spatial temperature signals at times $t$ and $t+\Delta t$, i.e. $\Delta\varphi$ For our purposes, Morlet was used as the mother wavelet and upon comparison with a Fourier based approach, both yielded similar results (see Fig. S4a in the Supplementary material). Specifically, a wavelet based approach allowed us to estimate the scale-dependent convection speeds separately for the positive and negative temperature structures. To do so, corresponding to a particular wavelet scale, the wavelet coefficients of the spatial temperature signals at a time $t$, were conditionally averaged only over those spatial locations where the temperature fluctuations were positive. An analogous operation was carried out for the negative counterparts and for a time $t+\Delta t$. We, therefore, defined the scale-dependent connection speeds of positive and negative structures as,
$U^{+}_{c}(\kappa)=\frac{\Delta\varphi_{\rm P}}{\kappa\times\Delta t},U^{-}_{c}(\kappa)=\frac{\Delta\varphi_{\rm N}}{\kappa\times\Delta t},$ (6)
In Eq. 6, the phase differences were only evaluated conditionally over the locations where temperature fluctuations were either positive or negative. This is denoted by the subscripts P and N, respectively.
@@ 3 Results and discussion
We begin with demonstrating the relationship between the convective speed estimations from the DTS measurements and the mean wind speed from the EC system at the clearcut. Corresponding to the EC data, $u$ and $w$ indicate the streamwise and vertical velocities, while $T,q,c$ indicate temperature, moisture, and $\rm CO_{2}$ concentrations. After employing a double-coordinate rotation, the turbulent fluctuations in these quantities are computed through linear detrending. By using both DTS and EC datasets, the impact of scale-dependent convection speeds on the turbulence statistics are discussed. We end by pointing out some potential limitations of the DTS measurements and how our findings influence in general the flux measurements from the eddy covariance systems.

### v2

@@ 2.3 DTS data processing
DTS data were post-field calibrated with Python (3.11.9) scripts and the processing followed des_tombe_estimation_2020, van_de_giesen_double-ended_2012 and peltola_probing_2022 with slight differences. Temperature can be estimated from the observed Stokes ($P_{s}$) and anti-Stokes ($P_{as}$) signals with the following equation (van_de_giesen_double-ended_2012):
$T_{x}(s,t)=\frac{\gamma}{\ln{\frac{P_{s}(s,t)}{P_{as}(s,t)}}+C(t)+\int_{0}^{s}\alpha(s^{\prime})ds^{\prime}},$ (1)
where $T_{x}=T_{x}(s,t)$ is the calibrated temperature at location $s$ along the fiber-optic cable at time step $t$, $\gamma$ and $C(t)$ are calibration parameters, and $\int_{0}^{s}\alpha(s^{\prime})ds^{\prime}$ describes the differential attenuation of $P_{s}$ and $P_{as}$ along the fiber. This integral was estimated from the forward and reverse signals for each location $s$ separately, corresponding to individual 30-min periods (van_de_giesen_double-ended_2012). Thereafter, at each time step, the calibration parameters $\gamma$ and $C$ were estimated for the forward and reverse signals by fitting Eq. 1 to $P_{s}$ and $P_{as}$ observed with the DTS instrument and $T$ measured with the reference sensors. Although, in theory, $\gamma$ should be constant, here it was treated as a fitting parameter varying in time. These parameters were used to calculate separate estimates of $T_{x}(s,t)$ from the forward ($T_{f}$) and reverse ($T_{r}$) signals. Since $T_{f}$ and $T_{r}$ were measured sequentially, $T_{r}$ data were linearly interpolated to match $T_{f}$ time stamps.
@@
$\sigma_{T}^{2}(s)=\frac{\overline{\left(T_{x}(s,t)-T_{x}(s+\Delta s,t)\right)^{2}}}{2},$ (2)
where $T_{x}$ represents either $T_{f}$ or $T_{r}$, $\overline{()}$ denotes temporal averaging and $\Delta s$ is the spatial step along the cable (0.127 cm). In this equation, we assume that the variability between consecutive data points in space is dominated by noise. The calibrated temperature $T$ at each $t$ and $s$ was calculated as a weighted average of $T_{f}$ and $T_{r}$ where inverse of noise estimates were used as weights (des_tombe_estimation_2020). Then the noise estimates for $T$ were calculated following noise propagation of weighted averaging. The spatial resolution of the data were reduced to 30 cm with spatial aggregation using weighted averaging as above and noise estimates for these spatial aggregates were calculated using noise propagation as above.
After this calibration procedure, we obtained calibrated $T$ at each time step $t$ and location $s$ along the cable. The $T$ data measured along the cable were converted to physical coordinates (ETRS89/TM35FIN coordinate system) using the measured cable locations (Sect. 2.2). Since there were double cabling at each physical location, the double measurements made at each location were further averaged using weighted averaging as above with the noise estimates obtained from the procedure described above. By utilizing this procedure, we obtained calibrated $T$ data with 2.3 s temporal and 30 cm spatial resolution throughout the measurement campaign at well-defined physical ($x$,$y$,$z$) locations (red lines in Fig. 1) over the study area. These calibrated temperature values were compared against the sonic anemometer measurements at the clearcut showing satisfactory agreement between the two (not shown).
@@ 2.4 Convection speed estimation
Figure 2: (a) The space-time contour plot of temperature fluctuations ($T^{\prime}$) is shown from the DTS measurements of a specific section (Section-1) for a particular 30-min period between 10:30-11:$0$ UTC on 18 May 2024. The spatial locations along the cable section are denoted by $X$ and the temperature fluctuations are obtained by linearly detrending the temperature signals measured along time. The white horizontal strips indicate the regions where the DTS cable was fastened to masts and the temperature values were omitted from those locations. (b) The space-time correlation curves of temperature fluctuations ($R_{TT}(\tau,r)$) are shown for the same 30-min period obtained from Section-1. The temporal lags or leads are denoted as $\tau$ while the colors indicate different values of spatial separations along $X$, $r$. The blue squares denote the peaks of the correlation curves ($\tau_{p}$) for each value of $r$. (c) A scatter plot between $r$ and $\tau_{p}$ is shown. The thick black color denotes the best fit line to compute the convective speed of the temperature structures, denoted as $U_{C}$.
@@
To use the DTS and EC datasets, we focus on buoyancy-dominated conditions when the temperature signals were sufficiently strong (see Fig. 10 of Appendix A for further discussion). In particular, we selected those 30-min periods when there was no rain, incoming shortwave radiation was larger than 300 $\rm W\ m^{-2}$ and it was not strongly varying within the 30-min periods, the sensible heat flux ($50\rm\ Wm^{-2}$) and variance of temperature ($0.1\rm\ K^{2}$) were sufficiently large, air temperature did not exhibit strong trends, and the direction of the mean wind was within $\pm 20^{\circ}$ of the DTS cable positions (see Fig. 1b). This wind direction criterion resulted in two different scenarios. For Sections 1 to 5, the mean wind was parallel to the forest edge while for Section 6 it was perpendicular to the edge (see Fig. 1b). After applying all these conditions, we ended up with a total of 35 hours of DTS and EC data, or in other words, 70 half-hour periods. The DTS dataset was used to estimate the convective speeds of the turbulent structures by using the temperature signal as a tracer. The EC dataset served as a testbed to evaluate the impact of these convective speeds on the routine flux measurements.
@@
Out of all our selected runs, for a specific representative 30-min period (10:30-11:00 UTC, 18 May 2024), we show the space-time contour plot of temperature fluctuations from Section 1 of the DTS cable (Fig. 2a). The temperature values were omitted from those locations where the DTS cable was fastened to wooden structures. This decision was undertaken to reduce any positive bias in the temperature signal since these wooden structures were not entirely thermally inert. Corresponding to each cable section, these omitted values represented around 6% of the cable length. Nevertheless, to ensure continuity in the data, those missing values were replaced with linearly-interpolated ones. Thereafter, the temperature fluctuations ($T^{\prime}$) were computed by linearly detrending the data over time separately at each location along the cable sections. The distance along Section 1 is denoted as $X$. To compute the convective speed of the temperature structures, we construct the space-time correlation curves from the DTS dataset. Mathematically, the two-point correlation functions of temperature fluctuations can be defined as,
@@
where $\tau$ are the temporal lags or leads, $r$ denotes the spatial separations along $X$, and $\sigma_{T(X)}$, $\sigma_{T(X+r)}$ are the standard deviations of temporal temperature fluctuations at locations $X$ and $X+r$. For the same data as shown in Fig. 2a, the quantity $R_{TT}(\tau,r)$ are plotted in Fig. 2b. Here, different colors represent the increasing values of $r$ while the blue squares indicate the peaks ($\tau_{p}$) attained by $R_{TT}(\tau,r)$ for a specific $r$. By fitting a straight line to a scatter plot between $r$ and $\tau_{p}$, the bulk convective speed of the temperature structures can be estimated (see the solid black line in Fig. 2c). The goodness of the fit can be adjudged by its $R^{2}$ value, which was well-beyond 0.9 in this example case (Fig. 2c). This linear relationship between $\tau_{p}$ and $r$ can be described as, $r=U_{C}\tau_{p}$, where $U_{C}$ is the bulk convective speed ($\rm m\ s^{-1}$). We restrict the $r$ values within 21 m while making the scatter plot, since for $r>21\rm\ m$, the $R_{TT}(\tau,r)$ values are small and therefore the estimations of $\tau_{p}$ are not statistically robust. To test the robustness of this method for atmospheric surface layer flows, we repeated this analysis on the cheng2017failure dataset (see Fig. S1 of the Supplementary Material). Similar to Figs. 2b–c, in Fig. S1 (Supplementary Material), the peak positions of $R_{TT}(\tau,r)$ curves shift towards increasing $\tau_{p}$ values as $r$ increases, thereby yielding a linear $r$-$\tau_{p}$ relationship. To ensure whether the shapes of $R_{TT}(\tau,r)$ curves were affected by the linear-interpolation operation, we increased the missing values to as large as 20% of the cable length and replotted Fig. 2b for this artificial data. We show this in Fig. S2 of the Supplementary Material and one could see that the peaks of $R_{TT}(\tau,r)$ curves behave analogously as in the original data.
Figure 3: (a) Contrary to Fig. 2b, the space-time correlation curves of temperature fluctuations ($R_{TT}(r,\tau)$) are shown where the $x$-axis represents the lags or leads in the spatial domain ($r$) while the colors indicate different values of temporal separations, $\tau$. The blue squares denote the peaks of the correlation curves ($r_{p}$) for each value of $\tau$. (b) A scatter plot between $r_{p}$ and $\tau$ is shown. The thick black color denotes the best fit line to compute the convective speed of the temperature structures, denoted as $U_{e}$.
Although the $r$-$\tau_{p}$ approach is more popular to compute the convective speed, everard2021sweeping argued that this method can give rise to high values due to the presence of sweeping effects associated with large turbulent eddies. Specifically, when the turbulence kinetic energy (TKE) of the flow is quite high, the sweeping effects have a strong domination on the estimates of $U_{C}$. To overcome this issue, everard2021sweeping suggested to use the $r_{p}$-$\tau$ relationship, which one can obtain by reversing the perspective, i.e. plotting the space-time correlation curves with respect to spatial lags or leads ($r$) at different values of temporal separations ($\tau$). For the same 30-min period akin to Fig. 2b, we plot the $R_{TT}(r,\tau)$ curves in Fig. 3a where the $x$-axis denotes the $r$ values while the colors of the curves indicate the values of $\tau$. The spatial peaks of $R_{TT}(r,\tau)$ curves are denoted as $r_{p}$ (see the blue squares in Fig. 3a). Similar to Fig. 2c, one can do a scatter plot between $r_{p}$ and $\tau$ (see Fig. 3b) to estimate a convection speed ($r_{p}=U_{e}\tau$), uncontaminated by the sweeping effects, which we denote as $U_{e}$. Upon comparing Fig. 2c with Fig. 3b, a difference emerges. In Fig. 3b, the $r_{p}$ values increase in a nearly consistent manner with $\tau$. On the other hand, in Fig. 2c, the $\tau_{p}$ values do not increase consistently but remain constant for certain ranges in $r$. This behavior occurs because the DTS data have a finer resolution in the spatial domain (30 cm) as compared to the temporal one (2.3 s).
After computing $U_{e}$, everard2021sweeping applied the elliptic model of he2010small to obtain the sweeping speed, $V$. The detailed derivation of the elliptic model of space-time correlation curves is provided in the Supplementary Material of everard2021sweeping, which we do not repeat here. By utilizing the relationship, $r=U_{C}\tau_{p}$, everard2021sweeping showed that the $V$ can be computed as (see their Eq. (5)),
$V=U_{e}\sqrt{\frac{U_{C}}{U_{e}}-1},$ (4)
where the non-zero values of $V$ describe any discrepancy between $U_{e}$ and $U_{C}$.
@@ 2.5 Scale-dependent convection speed
To compute the scale-dependent convection speeds of temperature structures, we follow the same procedure as outlined in cheng2017failure. By borrowing the concept of phase-dependent convection speeds from del2009estimation, cheng2017failure defined the convection speed at a wavenumber $\kappa$ ($U_{C}(\kappa)$) as,
$U_{C}(\kappa)=\frac{\Delta\varphi}{\kappa\times\Delta t}\$(5)
Here, $\Delta\varphi$ indicates the Fourier phase differences between the spatial temperature signals at time instants $t$ and $t+\Delta t$, $\kappa$ is the spatial wavenumber directly evaluated from the DTS measurements ($\kappa=2\pi/\lambda$, where $\lambda$ is the wavelength), and $\Delta t$ is the temporal lag (2.3 s for our DTS data). Unlike cheng2017failure, we used a wavelet based approach (torrence1998practical) to compute the phase differences between the spatial temperature signals at times $t$ and $t+\Delta t$, i.e. $\Delta\varphi$ . For our purposes, Morlet was used as the mother wavelet and upon comparison with a Fourier based approach, both yielded similar results (not shown).
@@ 3 Results and discussion
We begin with demonstrating the relationship between the convective speed estimations from the DTS measurements and the mean wind speed from the EC system at the clearcut. Corresponding to the EC data, $u$, $v$, and $w$ indicate the streamwise, cross-stream, and vertical velocities, while $T,q,c$ indicate temperature, moisture, and $\rm CO_{2}$ concentrations. After employing a double-coordinate rotation, the turbulent fluctuations in these quantities are computed through linear detrending. By using both DTS and EC datasets, the impact of scale-dependent convection speeds on the spectral properties of turbulence are discussed. We point out some potential limitations of the DTS measurements and how our findings influence in general the flux measurements from the eddy covariance systems.